# Quantitative surgery and total mean curvature

### v1

@@ 2. Total mean curvature bounds with fixed positive lower mean curvature bound
  • $H_{g_{i}}\geq\tfrac{\mu}{2}$ everywhere,
  • On $V_{i}$, the second fundamental form of $g_{i}$ is given by
    $\mathrm{II}_{g_{i}}=\tfrac{\mu}{2(n-1)}g_{M},$
@@
Let $\delta>0$ and let $\mathscr{C}:=\partial\Omega\times(-\delta,0]\subset\Omega$ be an open collar of $\partial\Omega$ in $\Omega$ induced by the normal exponential map along $\partial\Omega$ for $g_{i}$. Put $\mathscr{V}_{i}:=V_{i}\times(-\delta,0]\subset\mathscr{C}$. Equip the topological manifold
$\mathscr{V}_{i}\cup_{V_{i}\times 0}(V_{i}\times[0,1))=V_{i}\times(-\delta,1)\subset\Omega\cup_{M}W$
with the product smooth structure. Set $\widetilde{\Omega}_{i}:=\Omega\cup_{M}T_{i}\subset\Omega\cup_{M}W$. We have
@@
By construction, the smooth structures on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-{\delta},1))$ and on $\widetilde{\Omega}_{i}\cap{\rm int}(W)$ are compatible with each other. Hence, we obtain an induced smooth structure on $\widetilde{\Omega}_{i}$. We have $\partial\widetilde{\Omega}_{i}=\widetilde{M}_{i}$ as smooth manifolds. Furthermore, if $n\geq 7$, then $\widetilde{\Omega}_{i}$ is equipped with a spin structure inducing the one on $\widetilde{M}_{i}$.
We claim that there exists a metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ that satisfies
@@
    $H_{g_{\widetilde{\Omega}_{i}}}\geq\mu_{i}$ everywhere,
@@
    $\mathrm{scal}(g_{\widetilde{\Omega}_{i}})>{\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$.
Let us for now assume this metric $g_{\widetilde{\Omega}_{i}}$ exists and let $g_{\widetilde{M}_{i}}$ denote the induced metric on $\widetilde{M}_{i}$. Using $Z_{1}\cap Z_{2}=\emptyset$ and properties (i) and (ii), we obtain
(2.2) $\begin{split}\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq{}&\int_{\partial\Omega\setminus Z_{1}}H_{g_{\Omega}}{\rm dvol}_{g_{M}}+\int_{\partial\Omega\setminus Z_{2}}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\\ \leq{}&\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\widetilde{M}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\widetilde{M}_{2}}}\\ \leq{}&2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\widetilde{M}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\widetilde{M}_{2}}}\right\}.\end{split}$
By induction, for $i\in\{1,2\}$, Theorem 2.1 holds for the manifold $(\widetilde{M}_{i},\widetilde{g}_{i})$, the lower scalar curvature bound ${\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ and the lower mean curvature bound $\mu_{i}$. This concludes the induction step.
It remains to construct the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. For this, we first note that the restriction $h|_{V_{i}\times[0,1)}$ extends to a smooth metric $h^{\prime}$ on $V_{i}\times(-\delta,1)$ and that the $1$-jets of the metrics $g_{i}$ and $h^{\prime}$ along $V_{i}\times\{0\}\subset V_{i}\times(-\delta,1)$ coincide by the construction of $g_{i}$. Let $\chi\colon\partial\Omega\to[0,1]$ be a smooth cut-off function that is equal to $1$ on $U_{i}$ and equal to 0 on $\partial\Omega\setminus V_{i}$. Consider the smooth path of symmetric $(0,2)$-tensor fields $\psi\colon[0,1]\to C^{\infty}(T^{*}\mathscr{C}\otimes T^{*}\mathscr{C})$,
$\psi(t)(x,r):=(1-\chi(x)\cdot t)g_{i}+\chi(x)\cdot t\cdot h^{\prime}.$
Since $\chi(x)=0$ for $x\in\partial\Omega\setminus V_{i}$, this is well-defined. Possibly after passing to a smaller $\delta$ (hence to a smaller $\mathscr{C}$), the path $\psi$ consists of Riemannian metrics on $\mathscr{C}$, see Fig. 5.
@@
Figure 5. The local deformation $\psi$ for $t=1$.
By the choice of $\chi$, the $1$-jet of $\psi(t)$ along $\partial\Omega$ is independent of $t$. In particular, the scalar curvature of $\psi(t)$ along $\partial\Omega$ is given by
$\mathrm{scal}_{\psi(t)}=(1-\chi(x)\,t)\mathrm{scal}_{g_{i}}+\chi(x)\,t\cdot\mathrm{scal}_{h}>\min\{\sigma,\mathrm{scal}_{h}\}-1.$
The last lower bound is independent of the fill-in $(\Omega,g)$. By continuity, possibly after passing to a smaller $\delta$, which may depend on $(\Omega,g_{\Omega})$, we have, on $\mathscr{C}$,
$\mathrm{scal}_{\psi(t)} >\min\{\sigma,\mathrm{scal}_{h}\}-1,\quad t\in[0,1].$
We apply the local flexibility lemma [3, Theorem 1.2] to the local deformation $\psi$ on the open neighborhood $\mathscr{C}$ of $\partial\Omega$ in $\Omega$ and to the second order open partial differential relation of being a Riemannian metric with $\mathrm{scal}>\min\{C,\mathrm{scal}_{h}\}-1$ on symmetric $(0,2)$-tensor fields in $C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$. This results in a smooth path $\Psi\colon[0,1]\to C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$ of Riemannian metrics on $\Omega$ starting at $g_{i}$, and there exists some $0<\delta^{\prime}<\delta$, such that the restriction of $\Psi(t)$ to $\partial\Omega\times(-\delta^{\prime},0]$ coincides with $\psi(t)$, see Fig. 6.

### v2

@@ 2. Surgery for fill-ins with fixed lower mean curvature bounds
  • $H_{g_{i}}\geq\kappa$ everywhere,
  • On $V_{i}$, the second fundamental form of $g_{i}$ is given by $\mathrm{II}_{g_{i}}=\tfrac{\kappa}{n}g_{M}$,
@@
Let $\delta>0$ and let $\mathscr{C}_{i}:=\partial\Omega\times(-\delta,0]\subset\Omega$ be an open collar of $\partial\Omega$ in $\Omega$ induced by the normal exponential map along $\partial\Omega$ for $g_{i}$. Put $\mathscr{V}_{i}:=V_{i}\times(-\delta,0]\subset\mathscr{C}_{i}$. Equip the topological manifold
$\mathscr{V}_{i}\cup_{V_{i}\times 0}(V_{i}\times[0,1))=V_{i}\times(-\delta,1)\subset\Omega\cup W$
with the product smooth structure. We have
@@
By construction, the given smooth structures on mutual intersections are compatible with each other. Hence, we obtain an induced smooth structure on $\widetilde{\Omega}_{i}$.
Extend the restriction $h|_{V_{i}\times[0,1)}$ to a smooth metric $h_{i}^{\prime}$ on $V_{i}\times(-\delta,1]$ in an arbitrary way. The $1$-jets of the metrics $g_{i}$ and $h_{i}^{\prime}$ along $V_{i}\times\{0\}\subset V_{i}\times(-\delta,1)$ coincide by the construction of $g_{i}$. Let $\chi_{i}\colon\partial\Omega\to[0,1]$ be a smooth cut-off function that is equal to $1$ on $U_{i}$ and equal to 0 on $\partial\Omega\setminus V_{i}$. Consider the smooth path of symmetric $(0,2)$-tensor fields $\psi_{i}\colon[0,1]\to C^{\infty}(T^{*}\mathscr{C}_{i}\otimes T^{*}\mathscr{C}_{i})$,
$\psi_{i}(t)(x,r):=(1-\chi_{i}(x)\cdot t)g_{i}+\chi_{i}(x)\cdot t\cdot h_{i}^{\prime}.$
Since $\chi_{i}(x)=0$ for $x\in\partial\Omega\setminus V_{i}$, this is well-defined. Possibly after passing to a smaller $\delta$ (hence to a smaller $\mathscr{C}_{i}$), the path $\psi_{i}$ consists of Riemannian metrics on $\mathscr{C}_{i}$, see Fig. 6.
[…]
Figure 6. The local deformation $\psi_{i}$ for $t=1$.
By the choice of $\chi_{i}$, the $1$-jet of $\psi_{i}(t)$ along $\partial\Omega$ is independent of $t$. In particular, the scalar curvature of $\psi_{i}(t)$ along $\partial\Omega$ is given by
$\mathrm{scal}_{\psi_{i}(t)}=(1-\chi_{i}(x)\,t)\cdot\mathrm{scal}_{g_{i}}+\chi_{i}(x)\,t\cdot\mathrm{scal}_{h^{\prime}_{i}}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1.$
By continuity, possibly after passing to a smaller $\delta$, we have, on $\mathscr{C}_{i}$,
$\mathrm{scal}_{\psi_{i}(t)}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1,\quad t\in[0,1].$
We apply the local flexibility lemma [BH22, Theorem 1.2] to the local deformation $\psi_{i}$ on the open neighborhood $\mathscr{C}_{i}$ of $\partial\Omega$ in $\Omega$ and to the second order open partial differential relation of being a Riemannian metric with $\mathrm{scal}>\min\{\sigma,\min\mathrm{scal}_{h}\}-1$ on symmetric $(0,2)$-tensor fields in $C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$. This results in a smooth path $\Psi_{i}\colon[0,1]\to C^{\infty}(T^{*}\Omega\otimes T^{*}\Omega)$ of Riemannian metrics on $\Omega$ starting at $g_{i}$, and there exists some $0<\delta^{\prime}<\delta$, such that the restriction of $\Psi_{i}(t)$ to $\partial\Omega\times(-\delta^{\prime},0]$ coincides with $\psi_{i}(t)$, see Fig. 7.
[…]
Figure 7. The end of the global deformation $\Psi_{i}(1)$ obtained by applying the local flexibility lemma to the local deformation $\psi_{i}$.
The metric $\Psi_{i}(1)$ on $\widetilde{\Omega}_{i}\cap\Omega=\Omega$, the metric $h_{i}^{\prime}$ on $\widetilde{\Omega}_{i}\cap(U_{i}\times(-\delta^{\prime},1))$ and the metric $h$ on $\widetilde{\Omega}_{i}\cap W$ are compatible with each other. Hence, the union of these metrics defines a smooth metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. By construction, the following holds:
  • the metric on $\partial\widetilde{\Omega}_{i}$ induced by $g_{\widetilde{\Omega}_{i}}$ coincides with $g_{\widetilde{M}_{i}}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}=H_{g_{i}}=H_{g_{\Omega}}$ on $\Omega\setminus Z_{i}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}\geq\kappa$ on $Z_{i}\setminus U_{i}$,
  • $H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$ everywhere else on $\widetilde{M}_{i}$, by the choice of $T_{i}$,
  • $\mathrm{scal}_{g_{\widetilde{\Omega}_{i}}}>{\rm min}\{\sigma,\min\mathrm{scal}_{h}\}-1$.
Therefore, $g_{\widetilde{\Omega}_{i}}$ possesses all the required properties. This finishes the proof of Theorem (A). ∎
In preparation for Section 3, we prove the following result by a similar argument.
Lemma 2.1.
Assume that Theorem (C) holds in the case when $M$ is diffeomorphic to the $n$-sphere. Then Theorem (C) holds in general.
Proof.
Let $(M,g_{M})$ and $\sigma,\kappa\in\mathbb{R}$ be as in Theorem (C). Put $n:=\dim M$. As before, we can assume $n\geq 5$. Since $M$ carries only finitely many spin structures, we can assume without loss of generality that $M$ is equipped with a fixed spin structure and we are only considering spin fill-ins $\Omega$ of $M$ with isometries $\partial\Omega\approx M$ that are spin structure preserving. Let $B$ be a spin manifold with $\partial B=M$ such that $B$ induces the inverse spin structure on $M$.
Removing a small open disc from the interior of $B$, we obtain a compact spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$ and $\pi_{1}(W)$ by surgeries in the interior of $W$ (using $n\geq 5$), we can assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $1$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $2$.
Let $\ell\geq 0$ be the number of handles in this handle decomposition. The proof of Theorem (C) is by induction on $\ell$. The case $\ell=0$ is covered by the assumption of Lemma 2.1. Assume by induction that Theorem (C) holds for $\ell\geq 0$ and that the number of handles of $W$ is equal to $\ell+1$.
Let $(\Omega,g_{\Omega})$ be a spin fill-in of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$, $H_{g_{\Omega}}\geq\kappa$. Let $\omega\colon(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$ be a spin preserving Riemannian isometry. For simplicity of exposition, we will assume from now on that $\partial\Omega=M$ and $\omega=\operatorname{\mathrm{id}}$. This implies that $\Omega\cup_{M}B$ is a spin manifold.
We now proceed as in the proof of Theorem (A), but with one difference: For the construction of the surgery handles $T_{i}\subset W$ before (2.2), intersecting $M$ in $\mathrm{im}(\psi_{i})$, we do no longer refer to [LM84]. This implies that the lower bound $\widetilde{\kappa}$ for the mean curvatures of $\widetilde{M}_{i}$, $i=1,2$, in (2.2) may become negative. Note that, as before, $\widetilde{\kappa}$ does not depend on the fill-in $(\Omega,g_{\Omega})$ and that we can assume $\widetilde{\kappa}<\kappa$.
Set $\widetilde{\Omega}_{i}:=\Omega\cup T_{i}$. Note that $\widetilde{\Omega}_{i}$ carries an induced spin structure. By the construction below Eq. 2.3, which does not require $\widetilde{\kappa}>0$, we obtain metrics $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ with properties (i), (ii), (iii) and (iv). Using that $Z_{1}\cap Z_{2}=\emptyset$ and that $H_{g_{\Omega}},H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$, we hence obtain
(2.4) $\begin{split}&\int_{\partial\Omega}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}\\ &\qquad\leq{}\int_{\partial\Omega\setminus Z_{1}}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}+\int_{\partial\Omega\setminus Z_{2}}(H_{g_{\Omega}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\Omega}}\\ &\qquad\leq{}\int_{\partial\widetilde{\Omega}_{1}}(H_{g_{\widetilde{\Omega}_{1}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}(H_{g_{\widetilde{\Omega}_{2}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\\ &\qquad\leq{}2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}(H_{g_{\widetilde{\Omega}_{1}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}(H_{g_{\widetilde{\Omega}_{2}}}-\widetilde{\kappa}){\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}.\end{split}$
Here we take advantage of the fact that all integrands are non-negative.
Since $\widetilde{\kappa}$ and the metrics on $\partial\Omega$ and $\partial\widetilde{\Omega}_{i}$, $i=1,2$, are independent of $(\Omega,g_{\Omega})$, we obtain a constant $C$, which is independent of $(\Omega,g_{\Omega})$, such that
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq{}2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}+C.$
This concludes the induction step for Theorem (C). ∎
3. Spin fill-ins and creased initial data sets
This section concludes the proof of Theorem (C). According to Lemma 2.1, the only remaining case is when $M$ is diffeomorphic to the $n$-sphere. Let $(\Omega_{0},g_{0})$ be a spin fill-in of $(M,g_{M})$ such that
  • $\mathrm{scal}(g_{0})\geq\sigma_{0}$ for some $\sigma_{0}<0$,
  • $H_{g_{0}}\geq\kappa$ along $M$, w.r.t. the outward-pointing normal, for some $\kappa<0$.
In this section we adopt the sign-convention commonly used in general relativity, whereby the mean curvatures, denoted by $H(M\subseteq W)$, are computed with respect to the “infinity-pointing normal”, that is if $W\colon M\leadsto N$ is a cobordism from $M$ to $N$, then $H(M\subseteq W)$ is computed with respect to the interior unit normal, whereas $H(N\subseteq W)$ is computed with respect to the exterior unit normal.
Now, let $M,N$ be diffeomorphic to the sphere and consider the positive Lipschitz continuous function $h$ on $M$ given by
(3.1) $h_{0}=|\kappa|\sinh(1)+\frac{1}{3}H(M\subseteq\Omega_{0})_{+}>0.$
After possibly Replacing $h_{0}$ by a smooth approximation we may without loss of generality assume that $h_{0}$ is smooth. According to the proof of [SWW22, Lemma 3.1 and Theorem 4.1], there exists a Riemannian metric $g_{1}$ on $M\times[0,1]$ such that the cobordism $(\Omega_{1}=M\times[0,1],g_{1})$ connects $(M\approx S^{n},g_{M})$ to $N:=S^{n}_{r_{0}}$, equipped with the round metric $g_{\circ,{r_{0}}}$ of radius ${r_{0}}\gg 1$ independent of $h$, with the following properties:
@@ 3. Spin fill-ins and creased initial data sets
    $H(M\subseteq\Omega_{1})=h_{0}$ and $H(N\subseteq\Omega_{1})>0$ for the mean curvatures of $M,N$,
@@
    $\mathrm{scal}(g_{1})=\sigma_{1}$ for some constant $\sigma_{1}<0$ independent of the function $h_{0}$,
@@
    the total mean curvatures are related by $\int_{M}H(M\subseteq\Omega_{1})\leq\int_{N}H(N\subseteq\Omega_{1})$.
We note that the final inequality together with $H(M\subseteq\Omega_{1})\geq\tfrac{1}{3}H(M\subseteq\Omega_{0})$ implies
(3.2) $\int_{M}H(M\subseteq\Omega_{0})\leq 3\int_{N}H(N\subseteq\Omega_{2}).$
We define constants $c_{i}$, $i=0,1,2$ via
$c_{0}=\max\left\{4|\kappa|,\sqrt{\frac{n}{n+1}|\sigma_{0}|},(\cosh 1)^{-1}\sqrt{\frac{n}{n+1}|\sigma_{1}|}\right\},\qquad c_{1}=c_{2}=\cosh(1)c_{0},$
and we set
$\sigma_{2}=-c_{2}^{2}\frac{n+1}{n},$
which again is independent of $h_{0}$. We observe that
$c_{0}\geq 4|\kappa|,\qquad|\sigma_{2}|>\max\{|\sigma_{0}|,|\sigma_{1}|\},\qquad c_{i}^{2}\frac{n+1}{n}\geq-\sigma_{i}\qquad\text{for }i=0,1,2.$
As in the proof of [SWW22, Lemma 3.2 and Theorem 4.1], there exists an asymptotic hyperboloidal extension $(\Omega_{2},g_{2})$ of $N$ with $H(N\subseteq\Omega_{1})=H(N\subseteq\Omega_{2})$ and $\mathrm{scal}(g_{2})=\sigma_{2}$. Note that $\sigma_{2}=\sigma_{1}$ in [SWW22], while in our setting $\sigma_{2}<\sigma_{1}$ to compensate for the mean curvature jump at $M$. Consider now the manifold $(\Omega,g):=(\Omega_{0},g_{0})\cup_{M}(\Omega_{1},g_{1})\cup_{N}(\Omega_{2},g_{2})$ with the above choices for $h_{0}$ and $\sigma_{2}$ which has two corners at $M$ and $N$. The situation is illustrated in Fig. 8.
@@
Figure 8. At $M$ the mean curvature jumps from negative to positive. Treating $\Omega$ as an initial data set $(\Omega,g,k)$, this jump will have a favorable sign if $k$ is chosen appropriately, with an appropriate jump at $M$, which requires the scalar curvature of the asymptotic hyperboloidal extension to be sufficiently negative.
Similarly to [SWW22, Theorem 4.1 and p. 232–233] and [ST02], the functional
(3.3) $m(r)=\frac{1}{n\omega_{n}}\int_{\Sigma_{r}}\big(H(\Sigma_{r}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})-H(\Sigma_{r}\subseteq\Omega_{2})\big)\cosh(r\sqrt{-K})$
is monotonically increasing, and $m(r)\to m$ for $r\to\infty$. Here $K=\tfrac{\sigma_{2}}{n(n+1)}$ (=1=1 in [SWW22]), and $H(\Sigma_{r}\subseteq\mathbb{H}^{n+1}_{\sigma_{2}})=n\sqrt{-K}\coth(r\sqrt{-K})$ is the mean curvature of a sphere in hyperbolic space $\mathbb{H}^{n+1}_{\sigma_{2}}$ with constant scalar curvature $\sigma_{2}$. Moreover, $\Sigma_{r}$ denotes the $(r-r_{0})$-distance spheres to $N$ with respect to the hyperbolic background metric, e.g. we $\Sigma_{r_{0}}=N=S^{n}_{r_{0}}$. Finally, the hyperboloidal mass $m$ will be discussed in (3.4).
In view of (3.2) and (3.3), it suffices to prove $m\geq 0$ in order to establish Theorem (C). However, since $\kappa<0$ and $h_{0}>0$, there may be an unfavorable jump of the mean curvature at $M$, which corresponds to “infinitely negative” scalar curvature, as depicted in Fig. 8. This can be remedied by introducing a symmetric $(0,2)$-tensor $k$ on $\Omega$ which also jumps at $M$. With a suitable choice of $k$, the dominant energy condition for the resulting initial data set $(\Omega,g,k)$ is satisfied weakly across $M$. Hence, the hyperboloidal positive mass theorem can be applied in this more general context.
With the above choice of constants $c_{0},c_{1},c_{2}$, define the symmetric $(0,2)$-tensor $k_{i}=\tfrac{c_{i}}{n}g_{i}$ on $\Omega_{i}$ and let $k=k_{0}\cup k_{1}\cup k_{2}$. Consequently, we obtain an initial data set $(\Omega,g,k)$ satisfying the dominant energy condition
$\mu-|J|=\frac{1}{2}\left(\mathrm{scal}_{i}+c^{2}_{i}\frac{n+1}{n}\right)\geq\frac{1}{2}\left(\mathrm{scal}_{i}-\sigma_{i}\right)\geq 0\qquad\text{for $i=0,1,2$},$
where the energy density $\mu$ and momentum density $J$ are given by
$\mu=\frac{1}{2}\left(\mathrm{scal}+\operatorname{tr}_{g}(k)^{2}-|k|^{2}\right),\qquad J=\operatorname{div}(k-\operatorname{tr}_{g}(k)g).$
It remains to check whether the dominant energy condition of $(\Omega,g,k)$ holds weakly across $M$.
Given $\vartheta\in\mathbb{R}$, define the hyperbolic rotation matrix
$F_{\vartheta}=\begin{pmatrix}\cosh\vartheta&-\sinh\vartheta\\ -\sinh\vartheta&\cosh\vartheta\end{pmatrix}$.
Consider the spacetime mean curvature vectors
$\vec{H}(M\subseteq\Omega_{i})=\begin{pmatrix}H(M\subseteq\Omega_{i})\\ \operatorname{tr}_{M}(k_{i})\end{pmatrix}=\begin{pmatrix}H(M\subseteq\Omega_{i})\\ c_{i}\end{pmatrix}$.
Using the definition of $c_{1}$, we compute for $\vartheta=-1$
$F_{\vartheta}\vec{H}(M\subseteq\Omega_{0})-\vec{H}(M\subseteq\Omega_{1})$
$\qquad=\begin{pmatrix}\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}\\ \cosh(-1)c_{0}-\sinh(-1)H(M\subseteq\Omega_{0})\end{pmatrix}-\begin{pmatrix}H(M\subseteq\Omega_{1})\\ c_{1}\end{pmatrix}$
$\qquad=\begin{pmatrix}\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})\\ -\sinh(-1)H(M\subseteq\Omega_{0})\end{pmatrix}$.
Let $x\in M$ where $H(M\subseteq\Omega_{0})>0$. Using $c_{0}\geq 4|\kappa|$ and $2\sinh(1)>\cosh(1)$, we estimate in this case at $x$
$\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})$
$\qquad\geq\cosh(1)H(M\subseteq\Omega_{0})+(4-1)\sinh(1)|\kappa|$
$\qquad\geq\sinh(1)|H(M\subseteq\Omega_{0})|.$
Next, let $x\in M$ where $H(M\subseteq\Omega_{0})\geq 0$. Using that $\cosh(1)-\sinh(1)=e^{-1}>\tfrac{1}{3}$, we estimate in this case at $x$
$\cosh(-1)H(M\subseteq\Omega_{0})-\sinh(-1)c_{0}-H(M\subseteq\Omega_{1})$
$\qquad\geq\cosh(1)H(M\subseteq\Omega_{0})+(4-1)\sinh(1)|\kappa|-\frac{1}{3}H(M\subseteq\Omega_{0})$
$\qquad\geq\sinh(1)|H(M\subseteq\Omega_{0})|.$
This demonstrates that inequality (1.7) in [KKL25] holds at $M$, i.e. the dominant energy condition is weakly satisfied across $M$. We point out, using the notation of [KKL25], that $\nu_{+},\tau_{+}$ correspond to $\binom{1}{0},\binom{0}{1}$, and that the terms $\beta_{\pm}=k(\nu,\cdot)|_{TM}$, $\beta^{\Delta}=\beta_{+}-\beta_{-}-d\vartheta$ vanish, where $\nu$ denotes the infinity-pointing normal to $M$.
It remains to justify that the hyperboloidal positive mass theorem (PMT) can indeed now be applied, which is needed to resolve Gromov’s total mean curvature conjecture in the spin case. To this end, we first recall the standard hyperboloidal PMT for initial data sets. In what follows, let $b$ denote the standard $(n+1)$-dimensional hyperbolic metric with scalar curvature n(n+1)-n(n+1). We write $\mathcal{S}$ for the spinor bundle of the spin manifold $\Omega$, and set $\overline{\mathcal{S}}=\mathcal{S}\oplus\mathcal{S}$ for the corresponding spacetime spinor bundle, cf. [HZ25a, Definition 2.6], which admits a Clifford action of Minkowski space $\mathbb{R}^{n+1,1}=\operatorname{span}\{e_{0},e_{1},\dots,e_{n+1}\}$.
Theorem 3.1 ([CH04] and [HJZ25, Proposition 2.9]).